# for incompressible flows

### v1

@@ Abstract.
In these lecture notes, we provide a pedagogical introduction to the theory of mixing for incompressible flows from a PDE perspective. We discuss both the Lagrangian (ODE) and Eulerian (PDE, continuity equation) viewpoints, and introduce suitable notions of mixing scales that quantify the degree to which a scalar field transported by a velocity field becomes mixed. We then address the problem of establishing universal lower bounds on the time evolution of the mixing scale. This is first done in the smooth setting, using energy estimates and flow-based arguments, and later in the Sobolev setting, relying on quantitative estimates for regular Lagrangian flows. Finally, we present recent results concerning the sharpness of these lower bounds, their implications for the geometry and regularity of regular Lagrangian flows, and connections with more recent developments in the literature.
@@ 1. Introduction
which, under the divergence-free constraint (1.1) on the velocity field, is equivalent to the transport equation $\partial_{t}\varrho+u\cdot\nabla\varrho=0$. This equation models the evolution of a profile under the action of the prescribed velocity field. Physically, it expresses the local conservation of the mass of the advected quantity. We also recall that the two viewpoints (Lagrangian and Eulerian) are connected by the theory of characteristics, which provides a formula for the solution of (PDE) in terms of the flow:
@@
Exercise 1.1.
Prove formula (1.2).
@@
Show that the zero-average condition (1.3) is propagaged in time, that is, for every solution $\varrho$ of (PDE), it holds
@@
Figure 1. Numerical simulation of a mixing evolution (taken from [MR2801050]).
Our objective is to study the mixing of the passive scalar $\varrho$ solution of (PDE), namely its local convergence to zero (i.e., to its spatial average). Ideally, this occurs through stirring and the creation of fine filamentation, as illustrated in the numerical simulation in Figure 1 (taken from [MR2801050]). A natural idea would be to study the variance of $\varrho$ as a function of time. Thanks to the zero-average condition (1.3), the variance simply coincides with the square of the $L^{2}$-norm of $\varrho$. However, at least formally (that is, for sufficiently smooth solutions), every $L^{p}$-norm of $\varrho$ is conserved in time. In particular, the variance remains constant. Therefore, the study of the variance does not provide any useful information about the mixing of the passive scalar.
@@
Show that, for sufficiently smooth solutions of (PDE), any $L^{p}$-norm is constant in time. You can use the representation formula (1.2), or manipulate directly the first equation in (PDE).
@@
The variance (or the $L^{2}$-norm) is a suitable quantity to study mixing in the presence of diffusion, that is, for solutions to the advection-diffusion equation $\partial_{t}\varrho+{\rm div}\,(u\varrho)=\Delta\varrho$. In these notes we will not address this case.
Mixing occurs through filamentation and the intertwining of level sets of the passive scalar. A suitable way to capture this phenomenon is to consider weak convergence, which expresses the fact that local averages of $\varrho$ converge to zero (the spatial average of the passive scalar). Weak convergence to zero is the property that for every smooth test function $\varphi=\varphi(x)\in C^{\infty}(\mathbb{T}^{d})$,
@@
    Universal lower bounds. We establish universal lower bounds on the mixing scale, under natural energetic constraints on the velocity field. We do this via energy estimates (in Section 3), through Lagrangian estimates for Lipschitz velocity field (in Section 4), and by exploiting the theory of quantitative estimates for regular Lagrangian flows for Sobolev velocity fields from [MR2369485] (in Section 8).
@@ 2. The geometric mixing scale and the functional mixing scale
The mixing scale should be understood as the characteristic length at which the configuration appears approximately homogeneous. Put differently, if as observers we only resolve the system at a limited resolution, the mixing scale is the finest resolution at which the different phases of the fluid can no longer be distinguished. For instance, in Figure 3, the circles with solid boundary are too small: at that scale the two phases of the fluid remain clearly visible. The circles with dotted boundary have roughly the size of the mixing scale, since within each circle of such radius we see an essential balance between the two phases. Any circle with greater diamater, on the other hand, would be too large: the fluid looks mixed at that scale, but it is already mixed at finer scales as well. We would like to define the radius of the dotted circles as the mixing scale of the passive scalar.
@@ 2.1. The geometric mixing scale
Let $\{\varrho_{n}\}$ be equi-bounded in $L^{2}$ and with zero average. Show that:
@@ 2.2. The functional mixing scale
where the constant $c$ depends on the chosen normalization in the definition of the Fourier coefficients. Observe as a consequence that the zero-average condition (1.3) guarantees that (2.2) makes sense.
@@
Recall that the $L^{2}$-norm of a solution $\varrho(t,x)$ to (PDE) is conserved in time. Therefore, the fact that ${\rm mix}_{\rm f}(\varrho(t,\cdot))\to 0$ for a solution expresses that energy is being transferred from low to high frequencies. Indeed, for large values of $|k|$, the corresponding term in the sum in (2.2) contributes less to the total value, so the decay of the functional mixing scale reflects the progressive shift of energy toward higher frequencies.
@@ 2.3. A brief overview of fractional Sobolev spaces
If $s<0$, this makes sense if we assume $f$ to have zero average.
@@
Understand what could be a definition in Fourier variable of the homogeneous Sobolev space $\dot{W}^{s,p}(\mathbb{T}^{d})$, for $s\in\mathbb{R}$ and $1<p<\infty$.
If $f=f(x)$ is defined on $\mathbb{T}^{d}$ and $\nicefrac{{1}}{{\lambda}}\in\mathbb{N}$, then $f_{\lambda}(x)=f(\nicefrac{{x}}{{\lambda}})$ is also periodic, and hence well-defined as a function on $\mathbb{T}^{d}$. Denoting by $\nabla^{s}$ the derivative of order $s$, we formally have the relation
@@ 3. Energy estimates for the functional mixing scale
In this section we begin deriving lower bounds for the mixing scale. We follow [MR2801050] (see also [MR3026556]), where such bounds are obtained through energy estimates on the Eulerian side, namely for (PDE). These estimates are tailored to the functional mixing scale and have the advantage of being relatively straightforward and direct to derive. However, in certain cases they are not optimal. In what follows, we will consider three different types of energy budgets, one in each of the next subsections.
@@ 3.1. Velocity fields with uniformly-in-time Lipschitz regularity
${\rm Lip}\,(u(t,\cdot))\leq L\qquad\forall\,t\geq 0\,.$ (3.1)
Thanks to the zero-average condition (1.3), we can consider the unique $\varphi$ with zero average such that $\Delta\varphi=\varrho$ and multiply (PDE) by $\varphi$. We get
@@
$=-\int_{\mathbb{T}^{d}}\nabla\varphi\cdot\partial_{t}\nabla\varphi\,dx=-\frac{1}{2}\int_{\mathbb{T}^{d}}|\nabla\varphi|^{2}\,dx\,$,
@@ 3.2. Velocity fields with uniformly-in-time bounded kinetic energy
$\|u(t,\cdot)\|_{L^{2}(\mathbb{T}^{d})}\leq K\qquad\forall\,t\geq 0\,.$ (3.3)
We proceed as in Subsection 3.1, but we exploit a different rewriting for the second term:
@@
This, in principle, allows for $\|\varrho(t,\cdot)\|_{\dot{H}^{-1}(\mathbb{T}^{d})}=0$ and therefore $\varrho(t,\cdot)=0$ at some finite time. Such a phenomenon is called perfect mixing in finite time: the solution becomes completely mixed at that finite time. By time-reversal, perfect mixing in finite time also implies nonuniqueness for the initial-value problem for (PDE). In Section 5 we will see, through an example, that under a uniform-in-time bound on the kinetic energy as (3.3) both perfect mixing in finite time and nonuniqueness can indeed occur, and we will comment on the connections with the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory of flows of non-Lipschitz velocity fields.
@@ 3.3. Velocity fields with uniformly-in-time bounded enstrophy
$\|(u(t,\cdot))\|_{\dot{H}^{1}}\leq E\qquad\forall\,t\geq 0\,.$ (3.4)
In this case, we rewrite the second term as
@@
However, in dimension $d\geq 2$, estimate (3.5) does not hold. Even if it did, it would only yield a linear-in-time lower bound on the functional mixing scale (as in Subsection 3.2), which would not be sufficient to rule out perfect mixing in finite time and nonuniqueness for the initial-value problem for (PDE). Nevertheless, we will see in Section 8 that, differently from the situation in Subsection 3.2, under the assumption (3.4) uniqueness for the initial-value problem for (PDE) is guaranteed as a consequence of the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory, and therefore perfect mixing in finite time cannot occur. This shows that the lower bound (3.6) obtained through energy estimates cannot be optimal.
@@
Use Agmon’s inequality in three dimensions
$\|f\|_{L^{\infty}(\mathbb{T}^{3})}\lesssim\|\nabla f\|_{L^{2}(\mathbb{T}^{3})}^{\nicefrac{{1}}{{2}}}\|\Delta f\|_{L^{2}(\mathbb{T}^{3})}^{\nicefrac{{1}}{{2}}}$
and Brezis-Gallouët’s inequality in two dimensions
$\|f\|_{L^{\infty}(\mathbb{T}^{2})}\lesssim\|\nabla f\|_{L^{2}(\mathbb{T}^{2})}\left(1+\log\frac{\|\Delta f\|_{L^{2}(\mathbb{T}^{2})}}{\|\nabla f\|_{L^{2}(\mathbb{T}^{2})}}\right)^{\nicefrac{{1}}{{2}}}$
instead of the wrong estimate (3.5) and derive a lower bound (which necessarily allows for perfect mixing in finite time) on the functional mixing scale under the assumption (3.4). Heuristically, Agmon’s and Brezis-Gallouët’s inequalities express the fact that the $L^{\infty}$-norm can be controlled by the $L^{2}$-norm of the derivative, up to “paying with a bit of a higher-order derivative norm”.
@@ 4.1. Lower bound for the functional mixing scale
Recall the representation formula (1.2) for the solution of (PDE) in terms of the flow. If we now assume $\bar{\varrho}\in\dot{H}^{1}(\mathbb{T}^{d})$, we can prove an estimate for the propagation of the $\dot{H}^{1}$-regularity of the solution as follows:
@@
Provide an alternative proof of (4.5) by means of a direct energy estimate for (PDE), without relying on the representation formula (1.2).
@@
Even though the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory guarantees well-posedness of (PDE) below Lipschitz regularity, no analogue of the propagation of regularity (4.3), not even in a weaker form, holds in this regime, see the comments and references in Section 9. The interpolation inequality (4.4) is a true inequality: the two sides are not, in general, comparable.
@@ 4.2. Lower bound for the geometric mixing scale
We now present a direct proof of a lower bound for the geometric mixing scale, based on a purely Lagrangian argument. For simplicity, we restrict to the two-dimensional case and consider an initial configuration as in Figure 2 (left), with value +1+1 on the grey region and 1-1 on the white region. We pick the point $\bar{x}=(\nicefrac{{3}}{{4}},\nicefrac{{1}}{{2}})$ which is “in the middle” of the white region, and we denote ${\rm mix}_{\rm g}(\varrho(t,\cdot))=\varepsilon$.
@@ 5. Bressan’s mixing scheme and a glimpse of DiPerna-Lions-Ambrosio theory
To describe the evolution of the passive scalar, we first specify the configurations obtained at discrete times. We then show that there exists a velocity field for which these prescribed configurations coincide with the solutions of (PDE) at the given discrete times.
@@
Notice that, by time reversal, we have produced an example of nonuniqueness for the continuity equation (PDE). This was expected, based on the linear-in-time lower bound established in Subsection 3.2 under uniform-in-time integrability bounds on the velocity field. In fact, observing that $t\mapsto\frac{1}{1-t}$ is not integrable at $t=1$, this construction provides an instance of the sharpness of the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory (see also the review paper [MR3283066]). Indeed, this theory asserts (to keep things simple, for a bounded, divergence-free velocity field on the torus, even though these assumptions can be relaxed) that, assuming
$u\in L^{1}([0,T];W^{1,p}(\mathbb{T}^{d}))\;\;\text{for some $1\leq p\leq\infty$}\qquad\text{ or }\qquad u\in L^{1}([0,T];BV(\mathbb{T}^{d}))\,$,
then the continuity equation (PDE) admits a unique bounded weak solution, and there exists a unique suitable selection of solutions of (ODE) giving rise to the so-called regular Lagrangian flow associated with the velocity field. Moreover, stability holds in both cases. We will describe more refined examples showing the sharpness of these assumptions in Section 9. The DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory is based on the notion of renormalized solution and on the study of the convergence of suitable commutators, and it has been extended in [MR4071413] to the case of nearly incompressible $BV$ velocity fields, also answering in the positive the so-called Bressan’s compactness conjecture.

### v2

@@ Abstract.
In these lecture notes, we provide an introduction to the theory of mixing for incompressible flows from a PDE perspective. We discuss both the Lagrangian (ODE) and Eulerian (PDE, continuity equation) viewpoints, and introduce suitable notions of mixing scales that quantify the degree to which a scalar field transported by a velocity field becomes mixed. We then address the problem of establishing universal lower bounds on the time evolution of the mixing scale. This is first done in the smooth setting, using energy estimates and flow-based arguments, and later in the Sobolev setting, relying on quantitative estimates for regular Lagrangian flows. Finally, we present recent results concerning the sharpness of these lower bounds, their implications for the geometry and regularity of regular Lagrangian flows, and connections with more recent developments in the literature.
@@ 1. Introduction
which, under the divergence-free constraint (1.1) on the velocity field, is equivalent to the transport equation $\partial_{t}\varrho+u\cdot\nabla\varrho=0$. This equation models the evolution of a profile under the action of the prescribed velocity field. Physically, it expresses the local conservation of the mass of the advected quantity. We also recall that the two viewpoints (Lagrangian and Eulerian) are connected by the theory of characteristics, which provides a formula in terms of the flow:
@@
Exercise $1.$
Show that formula (1.2) provides a solution to (CE).
@@
Show that the zero-average condition (1.3) is propagaged in time, that is, for every solution $\varrho$ of (CE), it holds
@@
Our objective is to study the mixing of the passive scalar $\varrho$ solution of (CE), namely its convergence to zero (i.e., to its spatial average). Ideally, this occurs through stirring and the creation of fine filamentation. A natural idea would be to study the variance of $\varrho$ as a function of time. Thanks to the zero-average condition (1.3), the variance simply coincides with the square of the $L^{2}$-norm of $\varrho$. However, at least formally (that is, for sufficiently smooth solutions), every $L^{p}$-norm of $\varrho$ is conserved in time. In particular, the variance remains constant. Therefore, the study of the variance does not provide any useful information about the mixing of the passive scalar.
@@
Show that, for sufficiently smooth solutions of (CE), any $L^{p}$-norm is constant in time. You can use the representation formula (1.2), or manipulate directly the first equation in (CE).
@@
The variance (or the $L^{2}$-norm) is a suitable quantity to study mixing in the presence of diffusion, that is, for solutions to the advection-diffusion equation $\partial_{t}\varrho+{\rm div}\,(u\varrho)=\Delta\varrho$. In these notes we will not address this case and we refer the interested reader for instance to [MR4733112] and the references therein.
Mixing occurs through filamentation and the intertwining of level sets of the passive scalar: the scalar is stretched into thin, intertwined filaments by the flow, causing regions of different values to blend together more efficiently. We refer for instance to [MR3824693] for an introduction to the concepts of stirring, mixing, and transport. A suitable way to capture this phenomenon is to consider weak convergence, which expresses the fact that local averages of $\varrho$ converge to zero (the spatial average of the passive scalar). Weak convergence to zero is the property that for every smooth test function $\varphi=\varphi(x)\in C^{\infty}(\mathbb{T}^{d})$,
@@
    Universal lower bounds. We establish universal lower bounds on the mixing scale, under natural energetic constraints on the velocity fields. We do this via energy estimates (in Section 3), through Lagrangian estimates for Lipschitz velocity field (in Section 4), and by exploiting the theory of quantitative estimates for regular Lagrangian flows for Sobolev velocity fields from [MR2369485] (in Sections 8, after some preliminaries in Sections 6 and 7).
@@ 2. The geometric mixing scale and the functional mixing scale
The mixing scale should be understood as the characteristic length at which the configuration appears approximately homogeneous. Put differently, if as observers we only resolve the system at a limited resolution, the mixing scale is the finest resolution at which the different phases of the fluid can no longer be distinguished. For instance, in Figure 2, the circles with solid boundary are too small: at that scale the two phases of the fluid remain clearly visible. The circles with dotted boundary have roughly the size of the mixing scale, since within each circle of such radius we see an essential balance between the two phases. Any circle with greater diameter, on the other hand, would be too large: the fluid looks mixed at that scale, but it is already mixed at finer scales as well. We would like to define the radius of the dotted circles as the mixing scale of the passive scalar.
@@ 2.1. The geometric mixing scale
Let $\{\varrho_{n}\}$ be equi-bounded in $L^{\infty}$ and with zero average. Show that:
@@ 2.2. The functional mixing scale
where the constant $c$ depends on the chosen normalization in the definition of the Fourier coefficients. Observe as a consequence that the zero-average condition (1.3) guarantees that in (2.2) it does not really matter that we did not include the term $k=0$ in the sum.
@@
Recall that the $L^{2}$-norm of a solution $\varrho(t,x)$ to (CE) is conserved in time. Therefore, the fact that ${\rm mix}_{\rm f}(\varrho(t,\cdot))\to 0$ for a solution expresses that energy is being transferred from low to high frequencies. Indeed, for large values of $|k|$, the corresponding term in the sum in (2.2) contributes less to the total value, so the decay of the functional mixing scale reflects the progressive shift of energy toward higher frequencies.
@@ 2.3. A brief overview of fractional Sobolev spaces
Understand what could be a definition in Fourier variable of the homogeneous Sobolev space $\dot{W}^{s,p}(\mathbb{T}^{d})$, for $s\in\mathbb{R}$ and $1<p<\infty$. In fact, we will not really need the precise rigorous definition, but rather the operative properties of these spaces, so we will not focus too much on formal definitions.
If $f=f(x)$ is defined on $\mathbb{T}^{d}$ and $\nicefrac{{1}}{{\lambda}}\in\mathbb{N}$, then $f_{\lambda}(x)=f(\nicefrac{{x}}{{\lambda}})$ is also $1$-periodic, and hence well-defined as a function on $\mathbb{T}^{d}$. Denoting by $\nabla^{s}$ the derivative of order $s$, we formally have the relation
@@ 3. Energy estimates for the functional mixing scale
In this section we begin deriving lower bounds for the mixing scale. We follow [MR2801050] (see also [MR3026556]), where such bounds are obtained through energy estimates on the Eulerian side, namely for (CE). These estimates are tailored to the functional mixing scale and have the advantage of being relatively straightforward and direct to derive. However, in certain cases they are not optimal. In what follows, we will consider three different types of energy budgets, one in each of the next subsections.
@@ 3.1. Velocity fields with uniformly-in-time Lipschitz regularity
${\rm Lip}\,(u(t,\cdot))\leq L\qquad\forall\,t\geq 0$ (3.1)
for some constant $L$, where we denote by ${\rm Lip}\,(\cdot)$ the Lipschitz constant of a given function or vector field. Thanks to the zero-average condition (1.3), we can consider the unique $\varphi$ with zero average such that $\Delta\varphi=\varrho$ and multiply (CE) by $\varphi$. We get
@@
$=-\int_{\mathbb{T}^{d}}\nabla\varphi\cdot\partial_{t}\nabla\varphi\,dx=-\frac{1}{2}\frac{d}{dt}\int_{\mathbb{T}^{d}}|\nabla\varphi|^{2}\,dx\,$,
@@
Let us remark that assumptions on the Lipschitz regularity of the velocity field are usually too strong for applications in fluid dynamics; therefore, in the next two subsections we will weaken this assumption.
@@ 3.2. Velocity fields with uniformly-in-time bounded kinetic energy
$\|u(t,\cdot)\|_{L^{2}(\mathbb{T}^{d})}\leq K\qquad\forall\,t\geq 0$ (3.3)
for some constant $K$. We proceed as in Subsection 3.1, but we exploit a different rewriting for the second term:
@@
This, in principle, allows for $\|\varrho(t,\cdot)\|_{\dot{H}^{-1}(\mathbb{T}^{d})}=0$ and therefore $\varrho(t,\cdot)=0$ at some finite time. Such a phenomenon is called perfect mixing in finite time: the solution becomes completely mixed at that finite time. By time-reversal, perfect mixing in finite time also implies nonuniqueness for the initial-value problem for (CE). In Section 5 we will see, through an example, that under a uniform-in-time bound on the kinetic energy as (3.3) both perfect mixing in finite time and nonuniqueness can indeed occur, and we will comment on the connections with the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory of flows of non-Lipschitz velocity fields.
@@ 3.3. Velocity fields with uniformly-in-time bounded enstrophy
$\|u(t,\cdot)\|_{\dot{H}^{1}}\leq E\qquad\forall\,t\geq 0$ (3.4)
for some constant $E$. In this case, we rewrite the second term as
@@
However, in dimension $d\geq 2$, estimate (3.6) does not hold.
@@
Derive lower bounds on the functional mixing scale under the assumption (3.4). In two dimensions, this can be done by going back to (3.5) and using Gagliardo-Nirenberg’s inequality
$\|f\|_{L^{4}(\mathbb{T}^{2})}\lesssim\|\nabla f\|^{\nicefrac{{1}}{{2}}}_{L^{2}(\mathbb{T}^{2})}\|f\|^{\nicefrac{{1}}{{2}}}_{L^{2}(\mathbb{T}^{2})}\,$,
while in three dimensions one can use Agmon’s inequality
$\|f\|_{L^{\infty}(\mathbb{T}^{3})}\lesssim\|\nabla f\|_{L^{2}(\mathbb{T}^{3})}^{\nicefrac{{1}}{{2}}}\|\Delta f\|_{L^{2}(\mathbb{T}^{3})}^{\nicefrac{{1}}{{2}}}\,.$
In two dimensions, this gives the lower bound (3.7). However, (3.7) does not hold in higher dimension, at least not as a consequence of these energy estimates.
The lower bound (3.7) (valid in two dimensions) only yields a linear-in-time lower bound on the functional mixing scale (by the same computations as in Subsection 3.2), which is not sufficient to rule out perfect mixing in finite time and nonuniqueness for the initial-value problem for (CE). Nevertheless, we will see in Section 8 that, differently from the situation in Subsection 3.2, under the assumption (3.4) uniqueness for the initial-value problem for (CE) is guaranteed as a consequence of the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory, and therefore perfect mixing in finite time cannot occur. This shows that the lower bound (3.7) obtained through energy estimates should not be expected to be optimal.
@@ 4.1. Lower bound for the functional mixing scale
Recall the representation formula (1.2) for the solution of (CE) in terms of the flow. If we now assume $\bar{\varrho}\in\dot{H}^{1}(\mathbb{T}^{d})$, we can prove an estimate for the propagation of the $\dot{H}^{1}$-regularity of the solution as follows:
@@
Provide an alternative proof of (4.3) by means of a direct energy estimate for (CE), without relying on the representation formula (1.2).
Exercise 4.4.
Provide an alternative proof of (4.5) based on the duality formula
$\|f\|_{\dot{H}^{-1}(\mathbb{T}^{d})}=\sup_{\|\varphi\|_{\dot{H}^{1}(\mathbb{T}^{d})}=1}\int_{\mathbb{T}^{d}}f(x)\,\varphi(x)\,dx\,.$
In fact, this proof gives a better version of (4.5), since it does not require any regularity of $\bar{\varrho}$.
@@
Even though the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory guarantees well-posedness of (CE) below Lipschitz regularity, no analogue of the propagation of regularity (4.3), not even in a weaker form, holds in this regime, see the comments and references in Section 9. The interpolation inequality (4.4) is a true inequality: the two sides are not, in general, comparable.
@@ 4.2. Lower bound for the geometric mixing scale
We now present a direct proof of a lower bound for the geometric mixing scale, based on a purely Lagrangian argument. For simplicity, we restrict to the two-dimensional case and consider an initial configuration as in Figure 1 (left), with value +1+1 on the grey region and 1-1 on the white region. We pick the point $\bar{x}=(\nicefrac{{3}}{{4}},\nicefrac{{1}}{{2}})$ which is “in the middle” of the white region, and we denote ${\rm mix}_{\rm g}(\varrho(t,\cdot))=\varepsilon$ (recall the definition of the geometric mixing scale in Subsection 2.1).
@@ 5. Bressan’s mixing scheme and a glimpse of DiPerna-Lions-Ambrosio theory
To describe the evolution of the passive scalar, we first specify the configurations obtained at discrete times. We then show that there exists a velocity field for which these prescribed configurations coincide with the solutions of (CE) at the given discrete times.
@@
Notice that, by time reversal, we have produced an example of nonuniqueness for the continuity equation (CE). This was expected, based on the linear-in-time lower bound established in Subsection 3.2 under uniform-in-time integrability bounds on the velocity field. In fact, observing that $t\mapsto\frac{1}{1-t}$ is not integrable at $t=1$, this construction provides an instance of the sharpness of the DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory (see also the review paper [MR3283066]). Indeed, this theory asserts (to keep things simple, for a bounded, divergence-free velocity field on the torus, even though these assumptions can be relaxed) that, assuming
$u\in L^{1}([0,T];W^{1,p}(\mathbb{T}^{d}))\;\;\text{for some $1\leq p\leq\infty$}\quad\text{ or }\quad u\in L^{1}([0,T];BV(\mathbb{T}^{d}))\,$,
then the continuity equation (CE) admits a unique bounded weak solution, and there exists a unique suitable selection of solutions of (ODE) giving rise to the so-called regular Lagrangian flow associated with the velocity field. Moreover, stability holds in both cases. We will describe more refined examples showing the sharpness of these assumptions in Section 9. The DiPerna-Lions-Ambrosio [MR1022305, MR2096794] theory is based on the notion of renormalized solution and on the study of the convergence of suitable commutators, and it has been extended in [MR4071413] to the case of nearly incompressible $BV$ velocity fields, also answering in the positive the so-called Bressan’s compactness conjecture (this conjecture asserts the compactness of sequences of regular Lagrangian flows under incompressibility and $BV$ bounds, see [MR2033003]).